# The Alignment Trap: Complexity Barriers

### v1

@@ 5.2 Topological Scarcity Results
For a policy class with expressiveness $\text{EXP}(m)$, the fraction of policies that are perfectly safe is bounded by $2^{-2^{m}}$.
@@
5.3 Universal Impossibility Result
@@
For any countable set of alignment techniques $\{A_{1},A_{2},\ldots\}$, there exists a policy $\pi^{*}$ that is not aligned by any technique in the set.
@@
The proof uses a standard diagonalization argument. We construct a ”diagonal” policy $\pi^{*}$ that is specifically designed to misbehave in a way that the $i$-th alignment technique, $A_{i}$, fails to correct. By its construction, $\pi^{*}$ is misaligned after the application of any technique from the countable set. The full proof is in Appendix A.5. ∎
5.4 Invariance of Complexity under Alignment Transformations
Theorem 5.7 (Trap Universality).
For any alignment transformation that preserves the computational capability of a policy, the fundamental verification complexity barrier persists. Specifically, for any such transformation $A$, one can construct a policy $\pi$ with a cryptographic backdoor such that verifying the safety of the transformed policy $A(\pi)$ remains cryptographically hard (i.e., requires exponential time).
Proof Sketch.
The proof relies on constructing a policy $\pi$ whose catastrophic behavior is tied to a secret key of a pseudorandom function (PRF). An alignment transformation $A$ is assumed to be computationally efficient and ”capability-preserving,” meaning it cannot drastically simplify the policy’s logic. Such a transformation cannot learn the secret key from black-box access to $\pi$. Therefore, it cannot reliably remove the PRF-based trap without breaking the cryptography. The trap persists in the transformed policy $A(\pi)$, and verifying its safety remains as hard as breaking the PRF. The full proof is in Appendix E.4. ∎
5.5 Learning Impossibility Results
Theorem 5.8 (PAC-Bayes Alignment Lower Bound).
Let $\mathcal{H}$ be a hypothesis class and $S\subset\mathcal{H}$ be a set of perfectly safe policies. Assume $S$ has measure zero under any non-degenerate prior $P$ (i.e., $P(S)=0$). For any posterior distribution $Q$ with finite KL divergence from $P$, and any $\epsilon_{\min}>0$ representing the minimum catastrophic risk for policies outside $S$:
$\mathbb{E}_{h\sim Q}[L(h)]\geq\epsilon_{\min}\cdot(1-Q(S))$
where $L(h)$ denotes the catastrophic risk of policy $h$. If $\mathrm{KL}(Q\|P)<\infty$, then $Q(S)=0$, implying $\mathbb{E}_{h\sim Q}[L(h)]\geq\epsilon_{\min}>0$.
Proof Sketch.
(a) Counting Argument: There are $2^{2^{m}}$ possible Boolean functions on $m$ inputs. If safety corresponds to a unique function (e.g., the all-zero function), then only one of these $2^{2^{m}}$ functions is safe.
(b) Learning Argument (PAC-Bayes): The result from (a) implies that any non-degenerate prior over policies will assign a vanishingly small probability to the safe set, $P(S)\approx 0$. The PAC-Bayes inequality shows that to learn a posterior $Q$ that concentrates on the safe set (i.e., $Q(S)\approx 1$), the KL-divergence $\mathrm{KL}(Q\|P)$ must be infinite. This means a finite dataset cannot provide the information needed to overcome an unbiased prior and guarantee safety. The full proof is in Appendix E.2. ∎
5.6 Specification Instability Result
Theorem 5.9 (The Movable Goalpost / Epistemic Fragility Theorem).
Under conditions of stakeholder pluralism, where a set of stakeholders $L$ (potentially large or diverse) each define a set of catastrophic outcomes $B_{\text{cat}}^{(\lambda)}\subseteq Y$:
  1. 1.
    Aggregated Safety Requirement: A system $A$ is perfectly safe with respect to all stakeholders if and only if its outputs $A(x)$ avoid $B_{\text{cat}}^{\text{(total)}}=\bigcup_{\lambda\in L}B_{\text{cat}}^{(\lambda)}$ for all inputs $x$.
  2. 2.
    Emergent Impossibility from Pluralism (Core Statement): Even if each individual stakeholder’s catastrophic set $B_{\text{cat}}^{(\lambda)}$ is “reasonable” (e.g., $B_{\text{cat}}^{(\lambda)}\neq Y$, allowing for some non-catastrophic behaviors from their individual perspective, and perhaps even allowing for a non-zero error tolerance $\epsilon^{*}_{\lambda}>0$ with respect to their own ideal), the union $B_{\text{cat}}^{\text{(total)}}$ can become equivalent to the entire output space $Y$, or all but a measure-zero or computationally infeasible set of permissible behaviors. This forces the effective alignment error $\epsilon^{*}$ required to satisfy the collective to approach zero with respect to a target safe set $Y\setminus B_{\text{cat}}^{\text{(total)}}$ that may itself be vanishingly small or empty.
  3. 3.
    Epistemic Fragility: The overall safety definition $Y\setminus B_{\text{cat}}^{\text{(total)}}$ becomes epistemically fragile. The addition of a single new stakeholder with a sufficiently restrictive $B_{\text{cat}}^{(\text{new})}$ (or a change in an existing stakeholder’s definition) can drastically alter $B_{\text{cat}}^{\text{(total)}}$, potentially rendering a previously achievable collective safety target unachievable or an already safe system unsafe under the new aggregate definition.
6 Synthesis: The Alignment Trap
The preceding sections have established two independent barriers to ensuring AI safety: the Complexity Barrier, concerning the difficulty of finding and verifying safe policies, and the Specification Barrier, which addresses the challenge of defining safety itself. This section synthesizes these concepts to formalize the Alignment Trap, a dynamic that emerges from their interaction.
6.1 The Jaws of the Trap: Formalizing the Barriers
The two barriers are not merely conceptual; they are grounded in the formal results presented in the Core Theorems (Section 5) and are further elaborated here.
Remark 6.1 (Plain-English Summary of the Verification Barrier).
This theorem makes a very strong claim: for any AI system that could potentially cause a catastrophe, and for *any* safety mechanism or alignment technique we design to prevent such catastrophes, the problem of definitively checking whether this safety mechanism actually works perfectly (i.e., guarantees no catastrophes) is undecidable. This means there’s no universal algorithm that can always give a ”yes” or ”no” answer to whether our safety fix is foolproof. The difficulty lies in the very nature of detecting potential catastrophes.
Theorem 6.2 (The Verification Barrier).
For any system S with the catastrophic error property, and any alignment function F guaranteeing no catastrophic outputs, verifying F’s success is undecidable.
Remark 6.3 (Plain-English Summary of Measure-Zero Safety).
This theorem strengthens the ”needle in a haystack” argument. It suggests that even if we’re not trying to make an AI behave exactly like one specific ideal safe model, but are instead looking for *any* AI behavior that manages to avoid all known catastrophes, such universally safe behaviors are still exceptionally rare (measure-zero) in the space of all possible complex AI behaviors, especially if any small error can lead to disaster (Regime A).
Theorem 6.4 (Universal Measure-Zero Safety).
In Regime A (Effective Brittleness), the set of ALL functions (not just identity-preserving ones) that successfully prevent catastrophes has measure zero in a properly formalized function space.
6.2 The Vicious Cycle of Capability and Risk
The Alignment Trap emerges from a feedback loop driven by increasing AI capability. As a system’s capability ($C$) grows, so does its potential impact, leading to a societal demand for higher levels of safety. This demand translates into a requirement for a vanishingly small error tolerance ($\epsilon\to 0$).
This requirement for near-perfect safety activates both barriers simultaneously:
  • The demand for a low $\epsilon$ makes verification computationally intractable, as shown by the Complexity Barrier (Theorems 5.1, 5.2).
  • The need to define safety with such high precision runs into the Specification Barrier, as the definition becomes increasingly complex and contested (Theorem 5.9).
This creates a vicious cycle: the pursuit of greater capability leads to safety requirements that are impossible to either verify or specify.
6.3 The Alignment Trap Theorem
This dynamic is formally captured in the paper’s central theorem.
Remark 6.5 (Plain-English Summary of the Alignment Trap).
This is the central theorem of the paper. It states that as AI becomes more powerful, society demands near-perfect safety (almost zero mistakes). However, at the same time, the very act of checking or verifying that such a powerful AI is indeed perfectly safe becomes mathematically impossible or computationally too expensive to carry out. So, precisely when we need safety the most, we lose the ability to confirm it. This creates a ”trap”: the more capable the AI, the higher the safety standards, but the harder it is to meet and verify those standards.
Theorem 6.6 (The Alignment Trap).
As AI capability C increases:
  1. 1.
    Premise (Capability-Risk Scaling): Increasing capability $C$ leads to a requirement for alignment error $\epsilon_{required}(C)\to 0$.
  2. 2.
    Premise (Verification and Specification Barriers): Verifying that alignment error is below $\epsilon_{required}(C)$ is computationally intractable and defining safety with sufficient precision is non-computable or intractably complex.
  3. 3.
    Conclusion (The Trap): For any finite verification budget and specification complexity tolerance, there exists a capability threshold $C^{*}$ beyond which the required safety becomes impossible to achieve or verify.
6.4 Error Regimes and the Necessity of Perfection
The ”Fork Dichotomy” theorem provides additional context for why the demand for $\epsilon\to 0$ is often a technical necessity rather than a mere preference.
Theorem 6.7 (The Fork Dichotomy).
Any AI system, when evaluated against a concrete goal, must operate in one of two regimes:
  1. 1.
    Regime A (Hard Errors): Where any error, no matter how small, leads to a non-zero risk.
  2. 2.
    Regime B (Soft Errors): Where there exists an error buffer within which the system remains safe.
For systems operating in Regime A with respect to critical safety goals, the demand for zero risk is equivalent to a demand for zero error. This reinforces the Alignment Trap, as it connects the societal need for safety directly to the mathematical requirement for perfection that the core theorems show is intractable.
@@ 7.3 Paths Forward: The Strategic Trilemma
[…]
Figure 4: The two core theorems with quantifiable computational costs. Verification complexity shows operations required vs. target error rate for different system complexities $m$.
8 Conclusion
This paper has established that the pursuit of verifiably safe, highly capable AI systems confronts fundamental mathematical barriers. We have framed this challenge in terms of two interacting barriers, Complexity and Specification, which together create the Alignment Trap. Our results collectively demonstrate that verification becomes computationally infeasible long before it becomes practically necessary.
The convergence of these barriers forces a strategic trilemma: society must either constrain system complexity to maintain verifiable safety, accept unverifiable risks while scaling capabilities, or develop fundamentally new safety paradigms beyond verification. Our work provides the first systematic complexity-theoretic analysis of AI alignment and establishes rigorous bounds that any safety approach must confront. The path we choose will determine the future of our relationship with artificial intelligence.
@@ Appendix A Extended Proofs
Since $\mathsf{TAUTOLOGY}$ is $\mathsf{coNP}$-complete, $\varepsilon$-safety verification is $\mathsf{coNP}$-hard for any $\varepsilon<2^{-m}$. This establishes the $2^{-\Omega(m)}$ bound stated in the theorem. ∎ ∎
@@
Appendix B Proof of Theorem 5.2 ($\epsilon$-Robust Verification Complexity)
@@
Appendix C Additional Technical Results
C.1 Finite Verification Limitations
C.2 Empirical CRS Data Summary
@@
C.3 Proof of No Static Audit Completeness
@@ C.6 Inheritance conditions
C.7 $\varepsilon$-Bound Inheritance
Theorem C.1 ( $\varepsilon$-Bound Inheritance ).
@@ C.7 $\varepsilon$-Bound Inheritance
Appendix D Connection to Complexity Classes
@@
Our main theorem relies on the following foundational assumptions (which correspond to Axioms D.1, D.2, and D.3 in Section Additional Theorems and Proofs (D.22-D.26) of the main text):
Axiom D.1 (Geometric Thinness of $\Pi_{S}$).
@@
Axiom D.2 (Smoothness of Initial Condition Distribution).
@@
Axiom D.3 (Generic Dynamic Avoidance of Thin Sets).
@@
Remark D.4 (Justification of Axiom D.3).
Axiom D.3 encapsulates a key consequence of transversality theory (e.g., Thom’s Transversality Theorem) and dimensional arguments from geometric measure theory. Given Axiom D.1 ($\text{dim}_{\text{H}}(\Pi_{S})<n-1$), $\Pi_{S}$ has a codimension strictly greater than 1. A $C^{1}$ path $\phi_{\omega}$ traces a 1-dimensional curve. Generic $C^{1}$ paths (which occur for $\mu_{0}$-almost all $\omega$ under suitable genericity assumptions on the family $\{\phi_{\omega}\}$ and $\mu_{0}$) do not intersect sets of codimension $>1$. The dimensional mismatch ($\text{dim}(\text{Im}(\phi_{\omega}))+\text{dim}(\Pi_{S})-\text{dim}(\mathbb{R}^{n})<1+(n-1)-n=0$) implies an empty intersection for transverse cases.
@@
Lemma D.5 ($\Pi_{S}$ has Lebesgue Measure Zero).
Given Axiom D.1 and $n\geq 2$, the $n$-dimensional Lebesgue measure of $\Pi_{S}$ is zero: $\lambda^{n}(\Pi_{S})=0$.
Lemma D.6 (Initial Conditions Almost Surely Avoid $\Pi_{S}$).
Given Axiom D.1, Axiom D.2, and $n\geq 2$, the set of initial conditions $\omega$ for which $\phi_{\omega}(0)\in\Pi_{S}$ has $\mu_{0}$-measure zero:
@@
Theorem D.7 (Topological Alignment Trap).
Let $n\geq 2$. Under Axioms D.1, D.2, and D.3, the $\mu_{0}$-measure of the set of initial conditions $\omega$ for which the path $\phi_{\omega}$ ever intersects $\Pi_{S}$ is zero:
@@
Appendix E Impossibility of Universal Easy Verifiability for Capability-Preserving Alignment Transformations
Definition E.1 (Capability-Preserving Alignment).

### v2

@@ 5.4 Pillar IV: Information-Theoretic Impossibility (The “Can’t Store It” Barrier)
For a policy class with expressiveness capable of representing $N$ distinct behaviors, if only one of these behaviors is considered safe, the fraction of perfectly safe policies is $1/N$. For Boolean functions with $m$ inputs, this fraction is as small as $2^{-2^{m}}$.
@@
5.5 Pillar V: Dynamic Impossibility (The “Training Makes It Worse” Barrier)
This pillar establishes that the very process of making an AI more capable is actively hostile to safety.
Theorem 5.10 (The Capability-Safety Asymmetry).
In regions of high capability within the parameter space, the gradient for capability and the gradient for safety are generally anti-aligned. Consequently, optimizing for capability actively works against safety.
Proof Sketch.
Capability rewards generalizing broad statistical patterns, while safety requires adhering to sharp, arbitrary exceptions. The act of ”smoothing over” the data distribution to improve capability erodes the sharp boundaries that define safety, making a step that increases capability likely to decrease safety. ∎
Theorem 5.11 (Multi-Path Topological Alignment Trap).
Even with a polynomial number of training paths, the probability of finding a safe policy remains zero if the safe set is a lower-dimensional manifold.
Proof Sketch.
The proof relies on transversality theory. The probability of $n$ 1-dimensional paths intersecting a set of codimension greater than 1 remains zero for any $n$ that is polynomial in the space’s dimension. An exponential number of paths would be required for a non-zero chance of success. ∎
6 The Failure of Specification & Universal Solutions
Theorem 6.1 (The Movable Goalpost / Epistemic Fragility).
The definition of safety is unstable. As AI capabilities grow, or as more stakeholders are considered, the requirements for safety can expand, potentially rendering a previously safe system unsafe.
Proof Sketch.
The total set of unsafe behaviors is the union of unsafe sets defined by all stakeholders. Even if each stakeholder’s set is small, their union can cover the entire space of behaviors, making safety impossible. The addition of a single new stakeholder can render a previously safe system unsafe. ∎
Theorem 6.2 (Incompleteness of Static Audits).
For any finite audit suite of tests, there exists a policy that passes all tests but is catastrophically unsafe on an untested input.
Proof Sketch.
The proof relies on constructing a policy that behaves safely on the finite set of audited inputs but contains a ”backdoor” that triggers unsafe behavior on a specific, unaudited input. Since the set of all possible inputs is vastly larger than any feasible audit suite, such an untested input always exists. The full proof is in Appendix B.3. ∎
@@ 6 The Failure of Specification & Universal Solutions
For any countable set of alignment techniques, there exists a policy that is not aligned by any technique in the set.
@@
The proof uses a standard diagonalization argument. We construct a ”diagonal” policy that is specifically designed to fail for each technique in the list, ensuring it cannot be aligned by any of them. The full proof is in Appendix A.5. ∎
@@ 7.2 The $\epsilon>0$ Case: The ”Death Valley” of Safety
7.3 The Empirical & Practical Validation
The $\epsilon\leq 10^{-9}$ Requirement: For high-impact systems, society will demand near-perfect reliability (e.g., 9+ nines of safety). This low error tolerance makes the theoretical impossibilities practically relevant. Approximations are not good enough, and achieving this level of certainty is computationally, statistically, and physically impossible.
The Aviation Benchmark: Humanity’s best safety-critical system, commercial aviation, achieves $\sim$6 nines of safety after a century of learning from fatal crashes, with a massive infrastructure of restrictions in a highly controlled, physics-based domain. Demanding 9+ nines from AGI in an open, socially complex domain with no ability to learn from existential failure is a fantasy.
The Autonomous Vehicle Case Study: AVs are the perfect microcosm of this failure. They cannot even handle the ”easy” case of a single, well-defined domain because they are defeated at every level: they fail to handle enumerated traffic laws perfectly (Level 0), cannot handle unenumerated local conventions (Level 1), and are broken by unpredictable world-context interactions (Level 2). This demonstrates that even for a well-defined task like driving, with extensive real-world testing and billions in investment, we still can’t achieve safety because we can’t enumerate all the rules, handle world context, or predict all interactions.
7.4 Derivation of the $\epsilon\leq 10^{-9}$ Requirement
For high-impact systems, society will demand near-perfect reliability (e.g., 9+ nines of safety). This low error tolerance makes the theoretical impossibilities practically relevant. Approximations are not good enough, and achieving this level of certainty is computationally, statistically, and physically impossible.
If safety truly requires $\epsilon\leq 10^{-9}$:
  • This means a 1 in a billion error rate
  • One failure per billion operations
  • For an AI making 1000 decisions/second:
    • Expects catastrophic failure every $\sim$11.6 days
    • Over 1 year: $\sim$31 catastrophic failures
This connects directly to the paper’s core theorems because:
  • The verification complexity explodes at such low $\epsilon$
  • The measure of $910^{-}$ 9-safe policies is vanishingly small
  • No practical training can achieve this reliability
@@ 7.5 Paths Forward: The Strategic Trilemma
7.7 Final Synthesis: The Inescapable Conclusion
The findings of this paper point to a conclusion that is not about computational difficulty, but about fundamental structure. Our argument is centered on a core observation: We are using a technology (Machine Learning), whose premise is our inability to fully enumerate complex rules, to solve a problem (Safety) that requires adherence to every one of those non-enumerable rules. This establishes a circular dependency. The subsequent theorems presented are not independent difficulties but rigorous, mathematical confirmations of this central paradox. We have shown that from the perspectives of geometry, computational complexity, statistics, information theory, and system dynamics, this circle cannot be broken. Each line of formal inquiry converges on the same conclusion. The results stand on their own. Here are the standard safety engineering principles and how AI fails each one, proving there is no way to ”engineer” our way around the problem.
The findings of this paper can be synthesized by framing them not as abstract mathematical results, but as a direct violation of the foundational principles of safety engineering.
  1. Principle 1:
    The Hazard Analysis & Risk Assessment (e.g., FMEA, HAZOP)
    Standard Practice: Before building a bridge or a nuclear reactor, engineers must exhaustively enumerate all credible failure modes and their consequences. You must identify every way the system can fail (the hazard analysis) and then assess the risk. This enumeration is the absolute bedrock of safety.
    AI’s Violation: This is a direct collision with the Enumeration Paradox. The entire premise of using ML is that we cannot enumerate all the ”rules” of the real world. This also means we cannot possibly enumerate all the ways an AI can fail to follow those rules. An AI operating in an open world has an infinite, non-enumerable hazard space.
    The Killer Argument: ”You are asking to deploy a system for which you cannot even complete Step 1 of any standard safety certification process. You cannot list the hazards. Therefore, you cannot analyze the risk. This would be grounds for immediate project termination in any other engineering discipline.”
  2. Principle 2:
    The Requirement for a Complete Specification
    Standard Practice: To verify a system, you must have a complete, unambiguous specification of its required behavior. For an airplane, this is a multi-thousand-page document specifying exactly how it must behave under all specified flight conditions.
    AI’s Violation: This is the Specification Barrier. What is the complete specification for ”don’t manipulate humans” or ”respect cultural values”? It is infinite and unknowable. The goalpost is not just moving; it is in an infinite number of places at once.
    The Killer Argument: ”You are attempting to verify a system against a specification that does not, and cannot, exist. This is equivalent to asking an auditor to certify that a company complies with ’the law,’ without telling them which country’s laws to use.”
  3. Principle 3:
    The Principle of Validation & Verification (V&V)
    Standard Practice: V&V ensures that you (a) built the system right (verification) and (b) built the right system (validation). This involves testing the system against a finite, but comprehensive, set of test cases that represent the entire operational domain.
    AI’s Violation: This is where the Statistical Impossibility and Computational Impossibility pillars come in.
    Validation Failure: You cannot create a comprehensive test suite because of the Rare Disaster Paradox. The most critical test cases (e.g., preventing a novel pandemic) are impossible to generate.
    Verification Failure: Even if you had the tests, coNP-completeness proves that for a complex AI, you could not run the verification in the lifetime of the universe.
    The Killer Argument: ”Your proposed system is fundamentally invalidatable, as you cannot generate the most critical test cases. It is also unverifiable, as running the full test suite is computationally intractable. It fails both halves of the V&V requirement.”
  4. Principle 4:
    The Principle of Determinism and Predictability
    Standard Practice: Safety-critical systems must be deterministic or stochastically predictable. Given the same inputs and state, they must produce the same outputs, or a predictable distribution of outputs. This is essential for debugging and accident investigation.
    AI’s Violation: Complex AI models are stochastic black boxes. The sheer number of internal states and parameters makes them practically non-deterministic. Furthermore, the Capability-Safety Asymmetry proves that as they ”learn,” their failure modes change unpredictably. The system you test today is not the system you deploy tomorrow.
    The Killer Argument: ”You are deploying a system whose internal logic is opaque and whose failure modes actively shift during operation. In aerospace engineering, this would be called an ’unstable airframe.’ We don’t deploy unstable airframes and hope they learn to fly; we prove their stability on the ground first.”
The Synthesis: The Final, Devastating Conclusion
By using the language of standard safety engineering, we can frame the final argument like this:
”The development of advanced AI is not a bold new frontier of engineering. It is a wholesale abandonment of every single principle of safety that humanity has developed over the last 200 years of building things that can kill people.
We are building systems without knowing the hazards (Violation of Principle 1).
We are verifying them against specifications that don’t exist (Violation of Principle 2).
We are running tests that are fundamentally incomplete and computationally impossible (Violation of Principle 3).
And we are deploying them with the knowledge that their behavior is unstable and unpredictable (Violation of Principle 4).
Therefore, the problem is not that AI alignment is ’hard.’ The problem is that the entire endeavor, as currently practiced, is a form of engineering malpractice so profound that it would be criminally negligent in any other field. There is no ’getting around’ these principles; they are the definition of safety.”
7.8 Acknowledgements
@@ Appendix A Extended Proofs
Since $\mathsf{TAUTOLOGY}$ is $\mathsf{coNP}$-complete, $\varepsilon$-safety verification is $\mathsf{coNP}$-hard for any $\varepsilon<2^{-m}$. This establishes the $2^{-\Omega(m)}$ bound stated in the theorem. ∎
@@
A.6 Proof of Theorem 5.7 (PAC-Bayes Alignment Lower Bound)
Proof.
Assumptions:
  1. 1.
    The set of perfectly safe policies, $S$, has measure zero under a non-degenerate prior $P$, i.e., $P(S)=0$. This is justified by the scarcity results in Theorems 5.1, 5.2, and 5.9.
  2. 2.
The expected risk under the posterior distribution $Q$ is given by:
[…]
$\mathbb{E}_{h\sim Q}[L(h)]\geq\epsilon_{\min}$
A.7 Proof of Theorem 5.2 (Topological Alignment Trap)
A.7.1 Axioms for the Topological Alignment Trap
Axiom A.1 (Geometric Thinness of the Safe Policy Set).
$\text{dim}_{\text{H}}(\Pi_{S})<n-1$
Axiom A.2 (Smoothness of Initial Condition Distribution).
Axiom A.3 (Generic Dynamic Avoidance of Thin Sets).
A.7.2 Proof of the Theorem
The proof proceeds from these axioms:
  1. 1.
    From Axiom A.1, the safe set $\Pi_{S}$ has a Lebesgue measure of zero.
  2. 2.
    From Axiom A.2, the probability of initializing a policy within $\Pi_{S}$ is zero.
  3. 3.
    From Axiom A.3, a training path starting outside $\Pi_{S}$ will almost surely not intersect it.
Lemma A.4 ($\Pi_{S}$ has Lebesgue Measure Zero).
Given Axiom A.1 and $n\geq 2$, the $n$-dimensional Lebesgue measure of $\Pi_{S}$ is zero.
Proof.
Lemma A.5 (Initial Conditions Almost Surely Avoid $\Pi_{S}$).
Given Axioms A.1 and A.2, the set of initial conditions that lie within $\Pi_{S}$ has measure zero.
Proof.
Theorem A.6 (Topological Alignment Trap, Restated).
Under Axioms A.1, A.2, and A.3, the measure of the set of initial conditions for which the training path ever intersects $\Pi_{S}$ is zero.
Proof.
Let $E$ be the event that a path intersects $\Pi_{S}$. This can be decomposed into two disjoint events: starting in $\Pi_{S}$ ($E_{0}$) or starting outside and later entering $\Pi_{S}$ ($E_{+}$). From the preceding lemmas, the measure of $E_{0}$ is zero. From Axiom A.3, the measure of $E_{+}$ is also zero. Since the measure of the union of two measure-zero sets is zero, the total probability of intersection is zero. ∎
Appendix B Additional Technical Results
B.1 Proof of Theorem 5.4 ($\epsilon$-Robust Verification Complexity)
@@ Appendix B Additional Technical Results
B.2 Empirical CRS Data Summary
@@
B.3 Proof of No Static Audit Completeness
@@ B.6 Inheritance conditions
B.7 $\varepsilon$-Bound Inheritance
Theorem B.1 ( $\varepsilon$-Bound Inheritance ).
@@ B.7 $\varepsilon$-Bound Inheritance
Appendix C Connection to Complexity Classes
@@
Our main theorem relies on the following foundational assumptions (which correspond to Axioms C.1, C.2, and C.3 in Section Additional Theorems and Proofs (D.22-D.26) of the main text):
Axiom C.1 (Geometric Thinness of $\Pi_{S}$).
@@
Axiom C.2 (Smoothness of Initial Condition Distribution).
@@
Axiom C.3 (Generic Dynamic Avoidance of Thin Sets).
@@
Remark C.4 (Justification of Axiom C.3).
Axiom C.3 encapsulates a key consequence of transversality theory (e.g., Thom’s Transversality Theorem) and dimensional arguments from geometric measure theory. Given Axiom C.1 ($\text{dim}_{\text{H}}(\Pi_{S})<n-1$), $\Pi_{S}$ has a codimension strictly greater than 1. A $C^{1}$ path $\phi_{\omega}$ traces a 1-dimensional curve. Generic $C^{1}$ paths (which occur for $\mu_{0}$-almost all $\omega$ under suitable genericity assumptions on the family $\{\phi_{\omega}\}$ and $\mu_{0}$) do not intersect sets of codimension $>1$. The dimensional mismatch ($\text{dim}(\text{Im}(\phi_{\omega}))+\text{dim}(\Pi_{S})-\text{dim}(\mathbb{R}^{n})<1+(n-1)-n=0$) implies an empty intersection for transverse cases.
@@
Lemma C.5 ($\Pi_{S}$ has Lebesgue Measure Zero).
Given Axiom C.1 and $n\geq 2$, the $n$-dimensional Lebesgue measure of $\Pi_{S}$ is zero: $\lambda^{n}(\Pi_{S})=0$.
Lemma C.6 (Initial Conditions Almost Surely Avoid $\Pi_{S}$).
Given Axiom C.1, Axiom C.2, and $n\geq 2$, the set of initial conditions $\omega$ for which $\phi_{\omega}(0)\in\Pi_{S}$ has $\mu_{0}$-measure zero:
@@
Theorem C.7 (Topological Alignment Trap).
Let $n\geq 2$. Under Axioms C.1, C.2, and C.3, the $\mu_{0}$-measure of the set of initial conditions $\omega$ for which the path $\phi_{\omega}$ ever intersects $\Pi_{S}$ is zero:
@@
Appendix D Impossibility of Universal Easy Verifiability for Capability-Preserving Alignment Transformations
Definition D.1 (Capability-Preserving Alignment).